Lobachevskii Journal of Mathematics

http://ljm.ksu.ru
ISSN 1818-9962

Vol. 26, 2007, 137–137

\begin{matrix}c \\ ◯ \end{matrix} H. L. Huru

H. L. Huru

ERRATA: ADDENDUM TO “QUANTIZATIONS OF BRAIDED DERIVATIONS. 1. MONOIDAL CATAGORIES”,

LJM, VOL.XXXIV, AND “QUANTIZATIONS OF BRAIDED DERIVATIONS.

2. GRADED MODULES, 3. MODULES WITH ACTION BY A GROUP”, LJM, VOL.XXXV

In the paper ”Quantizations of braided derivations. 1. Monoidal categories”, in the Lobachevskii Journal of Mathematics, the following should be remarked: Proposition 1 can be removed and in Proposition 3 it is assumed that f and g are derivations.

In ”Quantizations of braided derivations. 2. Graded modules”, section 7.5 the results hold only for quantization of algebraic structure, not for the Poisson structure. The same goes for ”Quantizations of braided derivations. 3. Modules with action by a group”, section 6.5.